### v1

# Augmenting Smart Contract Decompiler Output through Fine-grained Dependency Analysis and LLM-facilitated Semantic Recovery

Abstract.
Decompiler is a specialized type of reverse engineering tool extensively employed in program analysis tasks, particularly in program comprehension and vulnerability detection. However, current Solidity smart contract decompilers face significant limitations in reconstructing the original source code. In particular, the bottleneck of SOTA decompilers lies in inaccurate method identification, incorrect variable type recovery, and missing contract attributes. These deficiencies hinder downstream tasks and understanding of the program logic. To address these challenges, we propose SmartHalo, a new framework that enhances decompiler output by combining static analysis (SA) and large language models (LLM). SmartHalo leverages the complementary strengths of SA’s accuracy in control and data flow analysis and LLM’s capability in semantic prediction. More specifically, SmartHalo constructs a new data structure - Dependency Graph (DG), to extract semantic dependencies via static analysis. Then, it takes DG to create prompts for LLM optimization. Finally, the correctness of LLM outputs is validated through symbolic execution and formal verification. Evaluation on a dataset consisting of 465 randomly selected smart contract methods shows that SmartHalo significantly improves the quality of the decompiled code, compared to SOTA decompilers (e.g., Gigahorse). Notably, integrating GPT-4o with SmartHalo further enhances its performance, achieving precision rates of 87.39% for method boundaries, 90.39% for variable types, and 80.65% for contract attributes.
Smart Contract, Decompilation, Static Analysis, Large Language Model
ccs: Software and its engineeringccs: Software and its engineering Software creation and managementccs: Software creation and management Software verification and validation
1. Introduction
Decompiler is a specific type of reverse engineering tool widely used for program analysis, such as program comprehension and vulnerability detection (Dinesh et al., 2020; Liao et al., 2024). For smart contracts written in Solidity, the goal of the decompiler is to recover the machine-executable code (e.g., the EVM bytecode (EVM, [n. d.])) back to the original source code written by developers (e.g., the Solidity program (Sol, [n. d.])). Unfortunately, in most cases, decompilers are incapable of fully reconstructing the original developer-written code, as the compiler irreversibly subverts crucial information for optimization (Hu et al., 2024). In particular, in Solidity EVM, critical information such as method boundaries, identifier names, and types are absent from the decompiler output. These limitations bring a number of negative impacts to its downstream tasks, as pointed out by quite a lot of prior research (Hu et al., 2024; Liao et al., 2022; Lai and Luo, 2020).
State-of-the-art solidity decompilers, such as Gigahorse (Grech et al., 2019) and Etherscan decompiler (eth, [n. d.]), mainly shares the following three deficiency in their output: (1) inaccurate method identification: current decompilers often fail to precisely determine method boundaries, leading to the omission of important methods or erroneous method range identification. For example, decompilers often incorrectly recover multiple methods as a single method, which obviously increases complexity in certain tasks (e.g., method-level similarity comparison). (2) inaccurate variable type recovery: decompilers may produce type errors that are inconsistent with static domain rules. For instance, the decompiler often ignores predefined type (e.g., the return value of hash function keccak256 is the type of bytes32), while recovering them as the type of uint256 uniformly. (3) lack of contract attributes. Smart contracts employ state variables to record critical contract attributes (e.g., Asset, identity, router). Although these contract attributes are explicitly stated in the source code through meaningful variable names and annotations, they are entirely omitted in the decompiled code. As highlighted in prior work (Liao et al., 2022), contract attribute is crucial for vulnerability detection tasks. In summary, the above deficiencies impede reverse engineers from understanding the program logic, and also adversely affect downstream tasks such as vulnerability detection and component analysis.
To the best of our knowledge, there has been very limited research (Liao et al., 2022; Chen et al., 2022b; Zhao et al., 2023) aimed at optimizing deficiencies for smart contract decompiler output. These research are far from satisfied, due to their limited scope and approach-wise weaknesses. More specifically, SmartDagger (Liao et al., 2022), a framework for detecting cross-contract vulnerability, can partially recover contract attributes (e.g., asset) for state variables from the decompiled bytecode. SigRec (Chen et al., 2022b) and DeepInfer (Zhao et al., 2023) can partially recover already-known method signatures and their corresponding parameter types. However, SigRec (Chen et al., 2022b) is heavily relied on pre-defined heuristics according to EVM instructions. Besides, both SmartDagger and DeepInfer are based on deep-learning models (e.g., deep neural networks) for information recovery. Therefore, their capability are quite limited to the model-scale and training datasets. For example, they suffers performance degradation when confronted with newly-emerging or rare contracts that are not exist in the training datasets.
Our work.  In this paper, we propose SmartHalo, a novel framework designed to optimize the output of existing smart contract decompilers. In particular, SmartHalo aims at recovering method boundaries, variable types, and contract attributes through a novel combination of static analysis (SA) and large language models (LLMs). The output of SmartHalo can facilitate a number of downstream tasks in program comprehension and security analysis. For example, eliminating false positives and false negatives for vulnerability detection.
The key insight of our research is that SA and LLMs can collaboratively enhance the output of existing decompilers. Specifically, the advantage of SA lies in its soundness, as it is accurate in handling the optimization targets that are with complex static constraints (He et al., 2018). Consider the optimization of variable types as an example. The types of state variables (e.g., address) are inherently related to the program expressions that access or modify them (e.g., msg.sender). Therefore, SA is well-suited to analyze these intricate static constraints (e.g., predefined functions) and refine the types of state variable accordingly. In the meantime, the strength of LLM lies in its completeness, as it possesses the flexibility to predict the optimization targets that lack static constraints (Chen et al., 2022a). For instance, with the remarkable capability of few-shot in-context learning, LLMs are flexible to predicting the types for certain local variables, which are rarely encountered in the program and often incapable of inferring their types based on the types of other variables.
SmartHalo extracts the static dependencies within the program through SA, and then leverages the superior generalization capabilities of LLMs to optimize the decompiled code after learning these static dependencies. This process involves three main steps, including dependency-based semantic extraction (Section 4.1), LLM-driven semantic enrichment (Section 4.2), and correctness verification (Section 4.3). In the first step, SmartHalo extracts three types of dependencies from the decompiled code, including state dependency (e.g., read and write on state variable), control-flow dependency, and type dependency, to construct a fine-grained Dependency Graph (DG). The DG contains static domain knowledge (i.e., dependencies) necessary for decompiler output optimization, which facilitates the LLMs in capturing them across the entire smart contract. In the second step, SmartHalo uses the DG to create prompts that include optimization target contexts (i.e., methods or variables), optimization output candidates (e.g., type or attribute), and chain-of-thoughts that represent inference steps of static analysis for optimization targets. By learning the prompts, LLMs enable the joint optimization on decompiler output with the advantages of both SA and LLMs.
A common challenge in LLM-adaptation is to handling its potential hallucination. In our task, LLMs unexpectedly alter the program behaviors of the original decompiled code, and introduce inference errors that contradict common static domain rules. As countermeasure, SmartHalo conducts the rigorous correctness verification for the LLM output. Specifically, SmartHalo utilizes symbolic execution and formal verification to validate the program-behavior equivalence between the original decompiled code and optimized code. In addition, SmartHalo integrates a set of static rules to identify and reject inference errors that violate static domain knowledge. To this end, the erroneous LLM outputs are combined with the verification suggestion to re-transmit to LLM for recurring optimization, for further eliminating the errors.
Evaluation.  To evaluate the effectiveness of SmartHalo, we randomly select 500 methods from the largest open-source contract dataset (xbl, [n. d.]), and manually labelled a dataset containing 456 pairs of source code and decompiler outputs of smart contract methods (44 methods encountered decompilation error). The evaluation results indicate that, compared to the original decompiler output (i.e., Gigahorse (Grech et al., 2019)), SmartHalo (with GPT-3.5) improves the precision of method boundary identification by 30.03% and enhances the precision of variable type inference by 12.01%. Additionally, SmartHalo successfully recovers 68.06% of contract attributes, significantly outperforming SmartDagger, which only recovers 10.07% of contract attributes. With the help of SmartHalo, 60.22% of optimized codes can directly be recompiled using the Solidity compiler, whereas the original decompiled output can not support recompilation. We also prove that the optimized output of SmartHalo enhances the effectiveness of downstream tasks, i.e., vulnerability detection. For example, SmartHalo enhances the precision by 21.96% and the recall by 38.00% on detecting integer overflow, improve the precision by 16.67% for contract attack identification. Particularly, we explore the impact of integrating GPT-4o with SmartHalo. With GPT-4o, SmartHalo further enhances its performance, achieving precision rates of 87.39% for method boundaries, 90.39% for variable types, and 80.65% for contract attributes.
In summary, this paper makes the following contributions:
  • We highlight the key limitations of current smart contract decompiler outputs that resistant various program analysis tasks in this domain.
  • We propose SmartHalo, a novel framework for optimizing smart contract decompiler output in a generic manner. We propose a set of novel mechanisms that combine static analysis and large language models, to establish accurate and flexiable optimization in decompiler output.
  • We perform extensive evaluation to show the effectiveness of SmartHalo. We demonstrate the efficacy of SmartHalo via three downstream tasks for vulnerability detection.
2. Background and Motivation
2.1. Smart Contract Decompiler
Decompilation plays a key role in program analysis by lifting bytecode back into a high-level representation. This process aims to elucidate the underlying logic of contract programs and facilitates a wide range of downstream tasks, such as vulnerability detection (Brent et al., 2020) and component analysis (Wu et al., 2023). Decompilation involves the recovery of variables, functions, and control-flow abstractions through various program analysis methods. Further, it then utilizes heuristic rules to synthesize these abstractions, thereby reconstructing the high-level code representation, commonly referred as decompiled code (Chen et al., 2022a). Recently, an investigation report reveals that over 99% of smart contracts do not make their source code publicly available, with only their bytecode being accessible (Con, [n. d.]). Therefore, smart contract decompilation is becoming increasingly important.
2.2. Deficiencies in Decompiler Output
The current decompiler output exhibits certain deficiencies that obstruct effective comprehension and analysis of the target program. Below we take smart contracts of Fig. 1, which contain source code and corresponding decompilation code, as the instances for illustrating the deficiency of decompiler output.
Deficiency-1. Inaccurate Variable Type Recovery.  Limited by the insufficiently generalized heuristic rule, the SOTA decompilers are inaccurate in recovering the variable type. This deficiency is evident in the example shown in Fig. 1(a). In the source code, the state variable uintStorage is a mapping type that maps bytes32 to uint256 (line 2). However, the decompiler incorrectly assigns it as a mapping uint256 to uint256 (line 10). Additionally, while decompilers narrow down the range of possible types for the parameter varg2 and assign it with bytes (line 11), the correct type of this parameter is actually a string array (line 3). Obviously, the SOTA decompilers commit errors in type inference. The notable efforts in this domain are SigRec (Chen et al., 2022b) and DeepInfer (Zhao et al., 2023), focus primarily on inferring function signatures and recovering parameter types within those signatures for smart contracts. However, since these tools concentrate on a subset of variable types, they are insufficient to address the broader deficiency of type recovery (e.g., variable uintStorage). In addition, the inaccuracies in variable type recovery result in significant difficulty in identifying specific vulnerabilities, such as overflow vulnerabilities (Lai and Luo, 2020).
Deficiency-2. Lack of Contract Attribute.  Due to the limited storage of blockchain, smart contract utilizes state variables (i.e., global variables) to record the key contract attributes (e.g., asset, identity, router). While contract attributes are crucial for the vulnerability analysis task, they are exceptionally difficult to be inferred from decompiled code (Liao et al., 2022). We take an example in Fig. 1(b) for illustration. Contract Bank exhibits a Reentrancy vulnerability because it records the asset balance after transferring assets. An adversary could exploit this by re-entering the contract to transfer assets without recording the change of asset balance (line 7-8). Thus, correctly identifying the state variable Acc.balance as an Asset attribute is critical for detecting this vulnerability. While the contract attribute is explicitly stated in the source code through meaningful variable names (i.e., Acc.balance), it is totally missed at the bytecode level and is labeled as store_0 in its decompiled code (line 11). Hence, inferring the contract attribute for the variable store_0 in the decompiled code becomes a formidable challenge. Without the contract attribute, it is difficult to pinpoint the vulnerability. Prior work (i.e., SmartDagger (Liao et al., 2022)) trains a neural machine learning model using a corpus of 1,200 smart contracts to recover contract attributes for state variables within smart contracts. However, due to limitations in model capacity and the size of the training set, SmartDagger degrade its accuracy of contract attribute prediction when encountering with newly-emerged or rare contracts.
Deficiency-3. Inaccurate Method Identification.  SOTA decompilers are notably ineffective in identifying methods within smart contracts due to their reliance on low-coverage heuristic rules. For instance, the Gigahorse (Grech et al., 2019) decompiler identifies methods by detecting call sites that are invoked recurrently. The heuristic rules result in poor performance when encountering complex methods, particularly inherited methods. In contrast to languages like C++ and Java, where inheritance is clearly delineated, smart contract inheritance involves embedding all base subcontracts ($B_{1},B_{2},...,B_{n}$) as code blocks directly into the inheriting contract $A$, without retaining explicit call information (Sol, [n. d.]). Hence, the heuristic rules failed in this case. As can be seen in Fig. 1(c), the inherited method validate is explicitly stated in the source code (line 6-8), but it is absent at the bytecode level. Our manual investigation reveals that the validate method is integrated into the handle method (lines 10-11). Consequently, SOTA decompilers like Gigahorse are unable to reconstruct such inherited methods, which are critical for downstream program analysis tasks such as cross-contract call flow analysis (Yang et al., 2024; Liao et al., 2022) and component analysis (Sun et al., 2023; Hu et al., 2017).
In conclusion, SOTA decompilers pose significant deficiencies, which have a negative impact on the different kinds of downstream program analysis tasks.
Figure 1. Three motivating examples for illustrating the limitations of current decompiler output.
2.3. Our solution
We propose our strategies for optimizing the smart contract decompiler output, combining with a discussion on potential solutions for the deficiencies presented in motivating examples of Fig. 1.
Strategy-1: Using SA to extract control-flow dependency, type dependency and static dependency.  First, type dependency is crucial for the variable type recovery. For example, the type error for variable uintStorage can be corrected by using SA to infer type dependencies, as illustrated in Fig. 1(b). Initially, the output of the predefined function keccak256 is of type bytes32 ($D_{t_{1}}$). Subsequently, the type of variable $v0$ depends on the return value type of the function _getTokenKey, which is also bytes32 ($D_{t_{2}}$). Finally, the key type of the variable setStorage depends on the type of $v0$ ($D_{t_{3}})$. Consequently, the type of setStorage should be corrected to mapping(bytes32= uint256). Second, control-flow dependency is critical for the method boundary identification. For instance, as shown in Fig. 1(a), while static constraints (i.e., call information) are unavailable in method validate, we can manually infer its boundaries by analyzing the contextual control flow, i.e., the method typically declares a new variable at the beginning (line 10) and returns the value at the end (line 11). Third, state dependency is important for contract attribute inference. We take the motivating example in Fig. 1(c) as an instance for illustration. By analyzing the dependency of the state variable, we can find that state variable stor_0 is written in line 16 ($D_{s_{1}}$) and line 18 ($D_{s_{2}}$). By manually analyzing their context usage, we can infer that they are utilized for asset transfer and asset receive, respectively. Hence, we can infer the attribute for stor_0 as the [Asset].
Strategy-2: Empowering LLMs to enrich semantics.  With the remarkable capability of few-shot in-context learning, LLMs are flexible to predicting the optimization target (i.e., variables or methods), which are rarely encountered in the program and often incapable of inferring them based on other variables or methods. We take Fig. 1(b) again as an instance for illustration. The type error for variable input can be corrected by using LLMs for prediction. Fig. 3 presents code optimized by LLM. Benefiting from the extensive training set and model capabilities, LLM can assign the variable input with string memory correctly (line 12). In addition, owing to its remarkable model capacity, LLMs are more effective and generalizable in learning static domain knowledge (i.e., dependencies) for optimization predictions, compared to SA’s manually crafting heuristic rules based on the same knowledge (Peng et al., 2023). For instances, instead of manually summarizing heuristic rules, we can utilize LLMs to learning the context control-flow for method boundary prediction, and to analyze the contextual usage of state variables to infer their contract attributes.
Inspired by the above analysis, our objective is to harness the advantages of both SA and LLMs to optimize decompiler output. Therefore, SmartHalo extracts three types of static dependencies including type dependency, state dependency and control flow dependency within the program through SA. Subsequently, SmartHalo leverages the superior generalization capabilities of LLMs to optimize the decompiled code after learning these static dependencies.
3. Design of SmartHalo
Figure 2. The decompiled code optimized by LLM for
the instance in Fig. 1(b).
Figure 3. An optimization error reported by LLM inference in terms of program-behavior non-equivalence.
To overcome the deficiencies, we propose a novel decompiler output optimization framework, called SmartHalo, based on hybrid SA and LLM. SmartHalo takes the decompiler output as input, and finally outputs the optimized code for analysts. Fig. 4 shows the overview of SmartHalo, which integrates the following mechanisms to ensure its effectiveness.
Dependency-based Semantic Extraction.  In this mechanism, SmartHalo aims to extract crucial semantics from the decompiled code through a fine-grain dependency analysis, which is then propagated into the LLM for further processing. To this end, our work complements and extends previous research by considering two new types of dependencies: namely, state dependency and type dependency. Specifically, SmartHalo analyzes three types of dependencies as the crucial program semantics, subsequently utilizing them to construct a Dependency Graph (DG). The extracted semantics (i.e., dependencies) represent the critical static domain knowledge necessary for decompiler output optimization, which facilitates the LLMs in capturing them across the entire smart contract (See Section 4.1).
LLM-driven Semantic Enrichment.  With the semantic re-presentation (i.e., DG) propagated by the above mechanism, SmartHalo constructs inference prompts to guide LLM to enrich the semantic of decompiler output. Specifically, according to the DG, SmartHalo utilizes our designed generator to constructs the context for optimization targets (i.e., methods or variables), for determining the code range related to optimization targets. Subsequently, SmartHalo further generates chain-of-thoughts that represent inference steps of static analysis for optimization targets, based on the dependencies among the DG, for teaching LLMs how to infer variable types, contract attributes and method boundaries. In addition, SmartHalo collects the candidates for variable types and contract attributes from the Solidity documentation (Sol, [n. d.]). Lastly, LLM learns from the above three types of prompt, and optimizes the crucial semantic (i.e., variable types, contract attributes and method boundaries) for decompiled code (See Section 4.2).
Correctness Verification.  To maintain the fidelity of the optimized code, SmartHalo should ensure that the optimized code preserves the consistent program behaviors with the original decompiler output and eliminate any inference errors that breach static domain rules.
However, verifying correctness is far from straightforward. First, due to the hallucination that commonly exist in LLMs, LLMs generates the random and unpredictable outputs that unexpectedly alter the program behaviors of the original decompiled code. For example, Fig. 3 illustrates a program behavior non-equivalence error caused by LLM optimization. In the original decompiled code, variables v0, v3 and v9 depends on parameters varg6, varg7 and varg8, respectively (line 02-04). However, in the optimized code, variables sigV, sigR and sigS incorrectly depend on the same parameter signaturesCount (line 06-08). Second, in addition to program behavior non-equivalence, LLMs also introduce inference errors that contradict common static domain rules, further complicating correctness verification. We take Fig. 3 again as an instance for illustrating such type of inference errors. In line 18, the value returned by the predefined function keccak256 is type of bytes32. According to Solidity documentation (Sol, [n. d.]), bytes32 cannot be directly converted into uint256, result in convertion errors.
SmartHalo implements two specific checks for establishing correctness verification. In the program-behavior equivalence check, SmartHalo utilizes symbolic execution and formal verification to verify the program-behavior equivalence between the original decompiled code and optimized code. In the static rule violation check, SmartHalo summarizes a set of static rules that are related to types. Note that static rule violation check donot consider method boundaries and contract attributes because there are no explicit static rules related to them in smart contract (Sol, [n. d.]). By applying these static rules, SmartHalo rejects the inference errors. (see Section 4.3)
Figure 4. The overview of SmartHalo.
4. Approach Details
4.1. Dependency-based Semantic Extraction
Contract Dependency Identification.  SmartHalo focus on identifying dependencies including the type dependency, state dependency, and control-flow dependency.
  • Type dependency. Type dependency refers to the dependency that the type of variable $v_{a}$ correlates to another variable $v_{b}$ or a specific expression $e_{b}$. Fig. 5 shows the syntax of all the expressions that generate types in smart contracts. Given the decompiled code, SmartHalo identifies the type dependency according to these syntaxes effectively.
  • State dependency. State dependency refers to the dependency between different state variables or dependency between state variables and expressions, such as read and write dependency on state variables. SmartHalo utilize SmartState (Liao et al., 2023), a SOTA state dependency analyzer to identify state dependency from given smart contracts.
Contract Dependency Graph Construction.  Subsequently, SmartHalo proposes the dependency graph (DG) as the representation of static domain knowledge for smart contracts.
The DG constructed by SmartHalo can be represented as a tripe $G_{c}=(N_{c},E_{c},X_{e})$. Specifically, SmartHalo encodes the following information: (1) the nodes of DG can be denoted as a set of all variables and related expressions in the decompiled code. Here, we use $N_{v}$ to represent the variable nodes, and leverage $N_{e}$ to represent the expression nodes. Hence, we have $N_{c}:=\left\{N_{v}\cup N_{e}\right\}$; (2) the edges of DG can be represented as a set of control-flow dependency edges $E_{d}$, state dependency edges $E_{s}$ and type dependency edges $E_{t}$. Similarly, we have $E_{c}:=\left\{E_{d}\cup E_{s}\cup E_{t}\right\}$. (3) $X(E_{c})\rightarrow\left\{DFD,SD,TD\right\}$ is a labeling function that maps an edge to one of the above three dependencies.
$e\in Expr::=$ $v~{}|~{}c~{}|~{}e~{}blop~{}e~{}|~{}e~{}numop~{}e~{}|~{}e~{}cmpop~{}e~{}|$
$e~{}bitop~{}e~{}|~{}(e,\dots,e)\ |~{}[e,\dots,e]~{}|~{}e(e,\dots,e)~{}|$
$e[e]~{}|~{}e[e:~{}e]~{}|~{}e.v$
Figure 5. The syntax of expression for typing in Solidity.
After generating control-flow graph, SmartHalo constructs the DG by incrementally adding state dependency edges and type dependency edges to the control-flow graph. Specifically, firstly, SmartHalo searches the whole control-flow graph to locate the variables and expressions that are related to type dependency. Further, SmartHalo utilizes SmartState to identify the variables and expressions that are related to state dependency. Finally, SmartHalo finds out the sources and targets among all these variables and expressions, and uses the direct edge to connect them in pair.
Figure 6. The optimization prompt for the instance in Figure 1 (b).
4.2. LLM-driven Semantic Enrichment
In this subsection, SmartHalo first generates domain-aware inference prompts according to the DG. Then, SmartHalo utilizes the inference prompts to guide the LLM to optimize the decompiler output. SmartHalo considers three types of inference prompts, including code contexts, inference candidates, and Chain-of-Thought (COT) prompts.
Code Context.  SmartHalo adapts different strategies to generate code snippets for variables and methods. The strategy for variables applies to optimizing variable type and contract attribute, and another strategy for methods is used to optimize method boundaries.
$\bullet$ Variables. To pinpoint the context related to target variables precisely, SmartHalo utilizes code slicing to identify code context fragments from the decompiled code for target variables, based on the DG. To this end, SmartHalo first generates the slicing DG based on the original DG. On the original DG, SmartHalo starts from target variable nodes and performs the forward and backward traversals (i.e., the same or opposite direction to DG edges). In this way, SmartHalo finds out all the nodes that have a dependency on target variables and generates a slicing subgraph of DG. With the slicing DG, SmartHalo combines the corresponding expressions together to generate a code snippet as the context prompt. In this way, SmartHalo narrows down the original decompiled code to the much smaller code snippet, which contains the information only related to decompiler output optimization of target variables.
$\bullet$ Methods. To generate the context prompt for target methods, SmartHalo searches the DG to find out the call chains that target methods lie in. Then, SmartHalo combines the corresponding methods in these call chains together to generate a code snippet as the context prompt.
We present the code context generation for the instance of Fig. 1(b), as highlighted in the gray part of Fig. 6. SmartHalo identifies all the expressions that have type dependency related to variable setStorage, and integrates them as the code context prompt.
Inference Candidate.  Subsequently, SmartHalo prompts the LLM with the candidates of variable type and contract attribute. In this way, SmartHalo can help the LLM narrow down the selection range to improve the accuracy.
Variable type candidates. Solidity is a statically typed language that provides a specific number of built-in types. In addition, the user-defined value type in Solidity is actually an alias that creates a zero cost abstraction over an elementary value type, so SmartHalo directly utilizes the elementary value type to represent it. According to Solidity documentation (Sol, [n. d.]), we collect the variable type candidates, as can be seen in the inference candidates (highlighted in green part) of Fig. 6.
Contract attribute candidates. In contrast to variable type, contract attribute is far from straightforward and should be summarized by representative smart contract dataset. We summarize the contract attribute candidates as a set [ Limit, Fee, Flag, Address, Asset, Router, Others ]  [^1] The details of manual summarizing process for contract attribute candidates can be assessed via Section Data Availability., and the detail of summarizing process are illustrated in Section 5.1
Chain-of-Thought Prompt.  COT prompts are actually a series of intermediate reasoning steps in static analysis. SmartHalo prompts the LLM with COT Prompt, for teaching LLM how to reason the correct method boundaries, variable type, and contract attribute from the perspective of static domain knowledge.
SmartHalo utilizes the slicing DG of the code slicing phase to generate the COT prompts. Given the slicing DG, SmartHalo organizes the nodes according to the distance between the nodes and the target variable node (i.e., hop counts). Specifically, the hop count of the target variable node is set as 0, and the hop counts of other nodes are calculated by their distance. SmartHalo translates each hop in the DG into a sentence of dependency description. Inspired by prior work (Peng et al., 2023), we summarize the effective COT prompt templates, as shown in Table 1. As can be seen, SmartHalo utilizes different COT prompt templates for the different optimization tasks. As can be seen, the category of type dependency is designed for the optimization on variable type recovery. Further, the category of state dependency is specific to the optimization on contract attribute recovery. The category of control flow dependency is used for the optimization on method boundary identification. To generate the COT prompt, SmartHalo fills the corresponding variable type, variable name, usages and statement, into the templates.
To illustrate the COT prompt generation, we present the COT prompt for the instance of Fig. 1(b), as highlighted in yellow part of Fig. 6. As discussed above, there are three type dependency edges (i.e., $D_{t_{1}}$, $D_{t_{2}}$, $D_{t_{3}}$) for the instance of Fig. 1(b). Among them, SmartHalo utilizes $expression\rightarrow variable$ template to describe $D_{t_{1}}$, and leverages $variable\rightarrow variable$ template to describe $D_{t_{2}}$ and $D_{t_{3}}$. The sentences of all the corresponding edges are connected together according to the order, to form the COT prompt.
Table 1. Chain-of-Thought Prompt Template for LLM-driven semantic enrichment
Category Type Template
Type dependency Variable $\to$ Variable
The type of variable[NAME] is consistent with the type of variable [NAME]
Type $\to$ Expression
The value of predefined function/operands[NAME] of expression [STATEMENT] is type of [TYPE]
Type $\to$ Variable The variable of [NAME] is assigned from [TYPE]
Expression $\to$ Variable
The type of variable [NAME] depends on expression [STATEMENT]
Variable $\to$ Expression
The operand(s)/target(s)/key(s)/value(s) of expression [STATEMENT] is/are [TYPE].
State dependency State variable $\to$ State variable
The attribute of state variable [NAME] is correlated to the attribute of state variable [NAME]
State variable $\to$ Expression
State variable [NAME] is written by Expression [STATEMENT]
Expression $\to$ State variable
State variable [NAME] is read by Expression [STATEMENT]
Control flow dependency
Return Value
Expression [STATEMENT] is used for return the value
Variable declaration Expression [STATEMENT] is used to declare the variable
Variable Usages The variable [Name] is used for [USAGES]
[NAME] indicates the name of variable, [STATEMENT] indicates the STATEMENT of expression nodes,
[TYPE] indicates the name of type nodes. [USAGES] indicates the usages for the variable.
4.3. Correctness Verification
Program-behavior Equivalence Check.  SmartHalo aims to check the program-behavior equivalence between methods of original decompiled code and methods of optimized decompiled code. We use $m$ and $m^{\prime}$ to represent the two versions of the same method in the original decompiled code and optimized decompiled code.
SmartHalo conducts the symbolic execution on methods $m$ and $m^{\prime}$ for generating symbolic summary $s$ and $s^{\prime}$. If methods $m$ and $m^{\prime}$ are equivalent in terms of program behaviors, their symbolic summary $s$ and $s^{\prime}$ are also equivalent in terms of functionality. An equivalence assertion is a first-order logic formula $\Phi$ that is used to assert such equivalence (Person et al., 2008).
(1) $\Phi=\neg(s\Leftrightarrow s^{{}^{\prime}})$
SmartHalo utilize formal verification method to analyze symbolic summaries $s$ and $s^{\prime}$, and determine whether equivalence assertion $\Phi$ is satisfied. If $\Phi$ is satisfied, method $m$ is non-equivalent to method $m^{\prime}$ in terms of program behaviors, otherwise, they are equivalent to each other in terms of program behaviors. Further, SmartHalo integrates the formal verification method that relies on SMT solver (i.e., Z3 solver (de Moura and Bjørner, 2008)), to find out satisfying assignments for equivalence assertion $\Phi$.
Note that since the symbolic summaries generally contain non-linear constraints, it is difficult for the Z3 solver to solve them. Such non-linear constraints commonly result in the solving failure. We eliminate most of the solving failure by utilizing methods proposed in ARDiff (Badihi et al., 2020).
Static Rule Violation Check.  SmartHalo traverse throughout the optimized code, then leverages violation rejection rules to identify and reject incorrect variable types predicted by the LLM.
Here, we summarize the violation rejection rules  [^2] More details about the violation rejection rules can be found in our project repository integrated by SmartHalo. Each violation rejection rule is composed of two parts, including specific premises (contents above the line) and conclusions (contents below the line). They are organized as follows.
(2) $\pi\vdash e:\theta$
where $\pi$ represents the context that contains lists that assign variable types to expression patterns. In this form, $e$ refers to the expression, and we use $e_{1},...,e_{n}$ to represent different expressions. $\theta$ is the variable types. We use $\theta_{1},...,\theta_{n}$ to represent different variable types. A rule in this form is called a judgment or assignment. Our goal is to get the context $\pi$ that assigns variable types to all the variables in code.
Finally, the incorrect optimization results predicted by LLM would be integrated with verification suggestions to form the violation information. SmartHalo retransmits the violation information into LLM to infer the correct method boundaries, variable types, and contract attributes anew. Further, the interaction can be iterated continually between LLM-driven semantic enrichment and correctness verification, until SmartHalo amends all the optimization errors identified by correctness verification, or a maximum iteration limit is reached.
5. Evaluation
Research Questions.  We summarize the following research questions for evaluating SmartHalo.
  1. RQ1.
    How effective is SmartHalo in terms of decompiler output optimization?
  2. RQ2.
    How does SmartHalo perform compared to other state-of-the-art mechanisms?
  3. RQ3.
    How effective are individual components of SmartHalo in terms of helping SmartHalo improve the precision of decompiler output optimization?
  4. RQ4.
    Can the output of SmartHalo help reverse engineers understand the smart contract bytecode in the real world?
5.1. Implementation and Evaluation Setup
Implementation.  We utilize ChatGPT (GPT-3.5) to support the LLM component of SmartHalo, due to the significantly expensive overhead of GPT fee. Additionally, we also conduct a small-scale experiments ( Section 5.2) to evaluate the effectiveness of SmartHalo that integrates the versions of GPT-4o, which shows SmartHalo indeed performs better with more advanced LLM models. Note that SmartHalo is designed to be model-agnostic and unlimited to a certain large language model, it offers the flexibility to integrate alternative large language models.
SmartHalo was implemented with around 1,799 lines of code in Python 3.8.10. All evaluation experiments were conducted on an Ubuntu 20.04 server equipped with an Intel i9-10980XE CPU (3.0GHz), an RTX 3090 GPU, and 250GB of RAM.
Evaluation Dataset and Ground-truth Establishment.  Previous research (Chen et al., 2022a; Lacomis et al., 2019) take significantly more effort to construct large datasets that encompass the different kinds of source code, as they are heavily dependent on the diversity of the training data. In contrast, one of the advantages of SmartHalo is its integration with LLMs (e.g., ChatGPT), which enables it to achieve strong performance without the need for extensive model training. Hence, we only require a relatively small testing dataset, rather than a large training dataset, to evaluate the effectiveness of SmartHalo.
Our evaluation dataset derives from a largest open-source smart contract dataset  (xbl, [n. d.]) as of Apirl 2024, totaling 963,151 smart contracts. From this original dataset, we randomly sampled 500 methods to construct the evaluation dataset. Ultimately, we filter out the methods the decompiler cannot handle, and obtain 456 pairs of source code and decompiler outputs of functions, for forming the final evaluation dataset. Here, we utilize method pairs instead of entire contracts because GPT integrated by SmartHalo requires shorter code snippets as input.
Furthermore, we established the ground truth for the evaluation dataset. Specifically, the ground truth for method boundaries and variable types was determined by manually comparing the decompiled code with the source code, as these aspects are explicitly stated in the source code. In contrast, the ground truth for contract attributes was established through manual inspection by our domain experts (i.e., researchers with three years of experience in smart contracts), who annotated the contract attributes via majority voting. As mentioned earlier, contract attributes refer to the semantic meaning represented by state variables (e.g., asset, address). To ensure comprehensive coverage of contract attribute categories, we also randomly sampled an additional 1,000 smart contracts from the original dataset and used a Large Language Model (LLM) to analyze the usage patterns of state variables within them. With the usage patterns, our experts manually summarize the representative categories as [ Limit, Fee, Flag, Address, Asset, Router, Others ].
Evaluation metrics.  We utilize four metrics for evaluation:
  • Method boundary match. We confirm the method boundary prediction is correct, by manually investigating whether the predicted method boundaries totally match the ground truth method boundaries.
  • Type match. Similar to DIRTY(Chen et al., 2022a), We confirm the type prediction is correct, if and only if the predicted types are totally matched the ground truth type. We serialize types to strings and determine the type matching by the string matching.
  • Contract attribute precision. We determine the precision of contract attributes by determining whether the predicted attributes are fully consistent with the ground truth attributes.
  • Recompilation failures. This metric represents errors yielded by the compiler when we are recompiling decompiled outputs. This indicates bugs or limitation in decompiler output (Williams-King et al., 2020).
5.2. Overall Effectiveness
To address RQ1, we evaluate the overall effectiveness of SmartHalo by comparing the optimized code generated by SmartHalo with the original decompiler output produced by the Dedaub decompiler (i.e., the web version of Gigahorse (Grech et al., 2019)) using four key metrics. To achieve this, we run SmartHalo on the evaluation dataset. Additionally, we conduct a detailed analysis of the reasons behind optimization failures. As can be seen in Table 2, the optimized code generated by SmartHalo significantly improves all four metrics compared to the original decompiler output. For example, the precision of method boundary identification is enhanced by 30.03%, and the precision of variable type inference is increased by 12.01%. Note that the the total number of variable types is inconsistent between Dedaub decompiler and SmartHalo, because the optimized code correctly recovers more methods which generate more parameter types. Moreover, while the original decompiler output lacks any contract attribute information, SmartHalo successfully recovers 68.06% of contract attributes. Additionally, unlike the original decompiler output, which is not suitable for recompilation, 60.22% of the optimized code generated by SmartHalo can be recompiled using a compiler.
Table 2. Comparison results between the optimized code generated by SmartHalo and original decompiler output produced by the Dedaub decompiler, for evaluate the overall effecctiveness of SmartHalo.
Method border
Variable type
Contract attribute
Recompilation failure
Original decompiler output (Baseline)
50.48% (316/626)
62.25% (676/1086)
N/A 100%
SmartHalo
80.51% (504/626)
84.26% (1349/1601)
68.06% (752/1102)
39.78% (185/465)
“NA” : the corresponding tool does not support optimizing this deficiency.
Table 3. Comparsion results between SmartHalo with GPT-3.5 and SmartHalo with GPT-4o.
Method border
Variable type
Contract attribute
SmartHalo with GPT-3.5
78.99% (94/119)
71.52% (113/158)
73.13% (98/134)
SmartHalo with GPT-4o
87.39% (104/119)
90.39% (254/281)
80.65% (100/124)
Optimization failures.  We manually inspect all the optimization failures of the four metrics. We found that most of optimization failures are caused by the limited model capability of GPT-3.5. As proved by our experiment in Table 3, SmartHalo equipped with GPT-4o can achieve precision rates of 87.39% for method boundaries, 90.39% for variable types, and 80.65% for contract attributes.
Exploration on GPT-4o.  Here, we explore how the model capability of LLM contributes to the overall effectiveness of SmartHalo. The study of the kind of LLM that can better drive SmartHalo is not the focus of our work. Conversely, we aim to explore how a more advanced LLM can further enhance the effectiveness of SmartHalo. GPT-4o is a more advanced LLM and performs better in various tasks compared with GPT-3.5. Hence, we compare the overall effectiveness in terms of three metrics between SmartHalo equipped with GPT-3.5 and SmartHalo equipped with GPT-4o. To save the expensive GPT fees, we randomly sample 50 method pairs from the evaluation dataset. We run the SmartHalo equipped with GPT-4o and SmartHalo equipped with GPT-3.5 overall the 50 method pairs. As can be seen in Table 3, SmartHalo equipped with GPT-4o further enhances the overall effectiveness. In conclusion, SmartHalo can integrate LLMs with better model capability to further enhance its effectiveness.
5.3. Comparison to Prior Work
To address RQ2, we compare SmartHalo with recent works. It’s important to note that we do not conduct comparative experiments regarding method boundary recovery, as there is a lack of existing work on optimizing method boundaries for smart contracts. In addition, SigRec (Chen et al., 2022b) is not open-sourced, and the model dataset for DeepInfer is unavailable for download (despite our inquiries with the authors, we received no response). Fortunately, we contacted the authors of SmartDagger (Liao et al., 2022) and obtained the artifact. Hence, we run SmartHalo and SmartDagger over the evaluation dataset to compare their precision in terms of contract attribute.
As can be seen in Table 4, SmartHalo significantly outperforms SmartDagger. For example, the precision of the contract attribute is improved by 44.69%. Further, we manually investigate the optimization failures generated by SmartDagger (Liao et al., 2022). Our findings indicate that the performance of SmartHalo degrades when encountering our dataset, which includes many newly-emerging smart contracts, due to limitations in its model capability and training dataset. Conversely, SmartHalo equipped with GPT-3.5 can maintain a higher precision of 68.06% for contract attributes. When equipped with GPT-4o, SmartHalo achieves even higher precision, with values of 80.65%.
Table 4. Comparison results between SmartHalo and SmartDagger in terms of contract attribute
Contract attribute
SmartDagger 10.07% (111/1102)
SmartHalo 68.06% (752/1102)
5.4. Ablation Study
As mentioned earlier, SmartHalo is composed of two components including static analysis and large language models. To answer RQ3, we compare the Dedaub decompiler, LLM (GPT-3.5) and SmartHalo in terms of method boundaries and variable types, to evaluate the contribution of two components to the overall effectiveness of SmartHalo. Note that we do not compare them in terms of contract attributes because LLM is incapable of outputting the contract attribute without our inference prompt (e.g., contract attribute candidate). To conduct this comparison, we further ran the LLM over the evaluation dataset. As can be seen in Table 5, compared with Dedaub decompiler, LLM presents a limited improvement in precision for method boundary identification but improves the precision of variable type inference by 10.01%. Note that the total number of variable types descends rapidly, because state variables and their types are unexpectedly deleted in the optimized code generated by LLM. We manually inspect the optimization failures of LLM. For method boundary identification, although LLM is capable of recovering a part of inherited methods, it also produces incorrect method boundary prediction which undermines the correct method boundaries in the original decompiler output due to the absence of static knowledge prompts (e.g., new variable declarations and return statements). Benefiting from the inference prompts generated by static analysis, SmartHalo improves the precision of method boundaries by 29.23% and enhances the precision of variable type by 12.52%, compared to LLM alone. In conclusion, both static analysis and the large language model significantly contribute to the improvement of decompiler output optimization in SmartHalo.
Table 5. Comparison results between SmartHalo and LLM.
Method border
Variable type
Original decompiler output (Baseline) 50.48% (316/626) 62.25% (676/1086)
LLM without static inference prompt
51.28% (321/626) 72.26% (375/519)
SmartHalo 80.51% (504/626) 84.78% (1349/1601)
5.5. User study
In this subsection, we highlight the practical effectiveness that SmartHalo helps reverse engineers comprehend the contract bytecode. We invite 13 domain experts and divide them into the professional group (i.e., 8 domain researchers with three years of experience) and intermediate group (i.e., 5 domain researchers with two years of experience), for evaluate the practical value of SmartHalo for different levels of reverse engineers. We randomly select 5 pairs of original decompiler output and optimized code of SmartHalo. Further, we invite each expert to review the code pairs and answer the questions listed in Table 6.
We calculated the average scores and variances for Q1-Q6 across the two groups, with the final results presented in Fig 8. The average scores range from 7.63 to 9.60 between the two groups, and both two groups exhibit relatively small variances within a range below 8. Given that a score of 5 represents a neutral attitude, this indicates that participants at different expertise levels demonstrated a relatively consistent positive attitude towards SmartHalo, particularly regarding method identification, type recovery, contract attributes, downstream tasks, program comprehension, and recompilation. Notably, the average scores from the professional group are lower than those from the intermediate group. To understand this, we interview the participants. The intermediate group views SmartHalo’s decompiler output optimization as comparable to that of a real-world domain expert, resulting in a more favorable response from them. However, due to the higher level of expertise, the professional group found the insights provided by SmartHalo to be less informative compared to the intermediate group. Additionally, the variances in the professional group were more stable than those in the intermediate group, which indicates that more experienced participants demonstrated more consistent and reliable recognition degree. Furthermore, we generated box plots to illustrate the score distribution in both groups, as shown in Fig.8. As can be seen, the scores of professional group are more well-distributed than those of intermediate group.
Table 6. The questions list in the user study.
Q1. Do you think the optimized code is more precise in terms of method borders?
Q2. Do you think the optimized code is more accurate in terms of variable types?
Q3. Do you think the contract attribute of optimized code is meaningful?
Q4. Do you think the optimized code is more helpful in comprehending bytecode?
Q5. Do you think the optimized code help the downstream program analysis tasks?
Q6. Do you think recompilation is an important improvement for decompiled code?
Figure 7. The average scores and variances.
Figure 8. The boxplot for the scores across two groups.
6. Effectiveness for Downstream Task
To evaluate the effectiveness of SmartHalo in supporting downstream program analysis tasks, we conducted three essential experiments. These experiments were designed to assess how effective the contract attributes, method boundaries, and variable types recovered by SmartHalo contribute to their respective downstream tasks.
Contract attributes for Reentrancy vulnerability detection.  As discussed earlier in Section 2.2, contract attribute is critical for locating certain types of vulnerability (e.g, Reentrancy). Prior research (Liao et al., 2022) shows that the effectiveness of contract attribute recovery reflects on eliminating the false positives for Reentrancy detection. Hence, We evaluate how the contract attributes optimized by SmartHalo contribute to improving the precision of Reentrancy detection. To this end, we utilize the latest manually-labeled DApp dataset introduced in previous research (Zheng et al., 2024), with a total of 81 positive labels for Reentrancy vulnerabilities. To conduct this evaluation, we run the SOTA tool SliSE (Wang et al., 2024) and another tool that integrates SliSE with SmartHalo over the manually-labeled DApp dataset to evaluate their precision. As can be seen in Table 6, SliSE+SmartHalo improve the precision to 80.41%, whereas the precision of SliSE is only 72.16%. We further manually inspect all the false positives reported by these two tools. We found that 8 of 27 false positives reported by SliSE can be eliminated by SmartHalo, with the help of more accurate contract attributes recovered by SliSE+SmartHalo. Fig. 9(a) illustrates a Reentrancy example of a false positive reported by SliSE. In this instance, SliSE erroneously flags the function recover due to its adherence to a pre-defined Reentrancy pattern, where function recover modifies state variables after transferring the asset (lines 8 and 9). Upon meticulous inspection of the source code, we found that the contract attribute of the state variable RecoveryDisable belong to [non-asset] (line 02). This attribute indicates that the recover cannot be exploited for profit gain by adversaries, thus recover does not pose a Reentrancy risk. In contrast, SmartHalo accurately identifies the contract attribute of this state variable as [Flag] (line 10), effectively eliminating this false positive.
Table 7. Comparison results for evaluating the effectiveness of SmartHalo on the Reentrancy detection
Reentrancy Precision
TP FP rate
SliSE 70 27 72.16%
SliSE+SmartHalo 78 19 80.41%
Table 8. Comparison results for evaluating the effectiveness of SmartHalo on the attack identification.
Attack contract identification Recall
TP FN rate
BlockWatchdog 15 3 83.33%
BlockWatchdog+ SmartHalo 18 0 100.00%
Table 9. Comparison results for evaluating the effectiveness of SmartHalo on Integer-overflow vulnerability.
Integer overflow vulnerability Precision Recall
TP FP rate TP FN rate
Mythril 13 5 72.22% 13 37 26.00%
Mythril+ SmartHalo 32 2 94.18% 32 18 64.00%
Method identification for attack identification.  Prior research (Yang et al., 2024) shows that accurate method identification is critical for locating certain types of tasks, such as attack identification. Specifically, attack identification relies heavily on the precise cross-contract call flow analysis between victim and attacker contracts, which effectively models the attack path. Incorrect method identification can lead to missed external calls and corresponding call flows, thereby impeding the identification of attack path. Therefore, the effectiveness of method boundary optimization is reflected in reducing false negatives in attack identification. To assess how method boundaries optimized by SmartHalo enhance the recall of attack identification, we utilize the latest attack contract datasets introduced by previous research (Yang et al., 2024), comprising 18 pairs of attacker and victim contracts from 15 real-world incidents. Further, we compare the SOTA tool BlockWatchdog (Yang et al., 2024) with an enhanced version that integrates SmartHalo with BlockWatchdog. As shown in Table 6, BlockWatchdog+SmartHalo enhances the recall by 16.67% compared with BlockWatchdog alone. Additionally, we manually inspect all the false negatives reported by BlockWatchdog, finding that BlockWatchdog+SmartHalo successfully eliminates all false positives through more accurate method identification optimized by SmartHalo. Fig. 9 (b) presents an example of false negative missed by BlockWatchdog. In this case, function LCtimeout contains a Reentrancy vulnerability, where it logs the transfer results (line 4) after transferring the asset (line 3). Further, an adversary could exploit this by invoking the function LCTimeout to trigger the Reentrancy vulnerability (path $A_{p1}$). BlockWatchdog fails to detect this attack path (i.e., false negative), because function LCtimeout is unavailable for BlockWatchdog at the decompiled code level (Yang et al., 2024). In contrast, SmartHalo accurately identify the method LCTimeout (line 13-15), thereby eliminating this false negative.
Variable types for Integer-overflow vulnerability detection.  Prior works demonstrate that variable type is crucial for locating certain types of overflow vulnerability, such as Integer overflow (Lai and Luo, 2020). For instance, detecting integer-overflow vulnerabilities heavily depends on accurately identifying the upper and lower bounds of variables based on their types. Incorrect variable type recovery results in both false positives and false negatives  (Lai and Luo, 2020). Therefore, the effectiveness of variable type optimization is reflected in its ability to reduce false negatives and false positives in integer-overflow vulnerability detection. We further evaluate the effectiveness of the variable types optimized by SmartHalo on the precision and recall of integer-overflow vulnerability detection. For this evaluation, we utilize integer overflow datasets collected from real-world attacks in previous research (Zhou et al., 2020), which includes 50 vulnerable contracts with integer overflow vulnerabilities. Further, we compare the SOTA tool Mythril (Consensys, [n. d.]) with a modified version that integrates Mythril with SmartHalo. As shown in Table 6, Mythril+SmartHalo enhances the precision by 21.96% and the recall by 38.00% compared with Mythril alone. Additionally, we manually investigate all the false negatives and false positives reported by Mythril. We found that 60% of the false positives and 51.35% of the false negatives reported by Mythril can be eliminated by Mythril+SmartHalo, thanks to the more accurate variable types optimized by SmartHalo. Fig. 9 shows an example of false negative missed by Mythril. In this case, the function batchTransfer contains a uint256 variable amount, which is correlated with two arguments, receivers and value, both of which can be passed arbitrarily via external calls (line 4), leading to an integer overflow vulnerability. Mythril fails to identify this vulnerability on the decompiled code level (i.e., false negative), as the decompiled code does not provide the type information for variable v3 (line 08). Conversely, SmartHalo correctly identifies the type of variable v3 as a uint256 type (line 7), thus eliminating the false negative.
Figure 9. The false alarms reported by SOTA tools, which can be actually eliminated by modified versions that integrate SOTA tools with SmartHalo.
7. Discussion and Limitation
SmartHalo shares the following advantages in decompiler output optimization: (1) As evidenced by evaluation, SmartHalo is obviously effective in optimizing method boundaries, variable types and contract attributes, demonstrating superior performance compared to previous approaches; (2) SmartHalo establishes a novel combination of static analysis (SA) and large language models (LLMs), enabling the joint optimization on the decompiler output with the advantages of both SA and LLMs, thereby overcoming the limitation of prior works; (3) SmartHalo implements a rigorous correctness verification for the optimization output to eliminate the optimization errors. With the above-illustrated advantages, SmartHalo can optimize decompiler output precisely and comprehensively. All of the developers, participants, and third-party authorities can utilize SmartHalo to facilitate a number of downstream tasks in program comprehension and vulnerability analysis.
Threats to validity.  An external threat to validity is that the number of smart contract methods used in our evaluation here is relatively small (i.e., 456 smart contract methods). However, establishing the evaluation dataset and its ground-truth requires a large manual effort and incurs expensive cost of GPT fees. In addition, the effectiveness of SmartHalo remains unaffected by the limitation of dataset size, because our manual investigation confirms that the dataset involves diverse and main application scenarios of smart contract [^3] Note that the manual investigation results can be available via Section Data Availability, such as decentralized governance voting, token implementation, cross-chain protocols, and gambling.
An internal threat to validity is the potential bias in annotating the ground-truth in the 456 smart contract method of the evaluation dataset. We have mitigated this threat by inviting three domain researchers performing the annotation independently, and annotated the method boundaries, variable types and contract attributes via a manner of majority voting. Another internal threat to validity is the predefined static rules employed by SmartHalo for correctness verification (Section 4.3). However, SmartHalo remains a highly generalizable framework for correctness verification, because the predefined static rules are derived from the syntax of smart contracts, as outlined in Solidity documentation (Sol, [n. d.]). Hence, these static rules are essential and universally applicable across all Solidity contracts. Note that SmartHalo is incapable of supporting analyzing non-Solidity smart contract, because the decompiled code generated by its core decompiler Gigahorse (Grech et al., 2019) is specifically written in Solidity. However, SmartHalo offers the flexibility to integrate alternative decompilers tailored to other languages to extend its capabilities in supporting smart contracts written in other languages. Importantly, the effectiveness of SmartHalo is still guaranteed by the fact that most smart contracts are written in Solidity language.
8. Related Work
Smart Contract Decompilation.  The decompilation of smart contracts has garnered significant attention due to the prevalence of bytecode-only smart contracts (Suiche, [n. d.]; Brent et al., 2018; Liu et al., 2023). Gigahorse (Grech et al., 2019) translates smart contracts from low-level EVM bytecode into a 3-address code representation. Elipmoc (Grech et al., 2022) employs a novel context sensitivity called transactional sensitivity to achieve a more effective static abstraction. Erays (Zhou et al., 2018) and EtherSolve (Contro et al., 2021) aim to recover the CFG from the EVM bytecode. Recently, SigRec (Chen et al., 2022b) has been proposed to recover the signatures of public functions. DeepInfer (Zhao et al., 2023) leverages deep learning techniques to infer function signatures.
Decompilation Optimization.  The existing decompilation optimization research mainly focus on other languages such as C++ and Java (Pei et al., 2020), which can be divided into the following two categories (Pang et al., 2021; Haq and Caballero, 2021). The first category aims at recovering the semantic information from the assembly code or intermediate language. Among them, previous works, including DEBIN (He et al., 2018), TIE (Lee et al., 2011), OSPREY (Zhang et al., 2021), and BDA (Zhang et al., 2019), analyze the program dependency via the static analysis. However, they are with low coverage due to their reliance on heuristic rules. Another group of works, including Nero (David et al., 2020), NFRE (Gao et al., 2021), SYMLM (Jin et al., 2022), focus on designing encoder-decoder architecture models to predict the method name for the binary. The other category of work focuses on optimizing variable name, variable type, and structure of code generated by decompiler. For example, Jaffe et al. utilize statistic machine translation to recover variable names (Jaffe et al., 2018). DIRE (Lacomis et al., 2019) leverages neural machine translation to predict variable names. DIRTY (Chen et al., 2022a) with the transformer-based model to predict the variable types and names. DeGPT (Hu et al., 2024) applies the LLM to augment the decompiler output. However, these works ignore the static domain knowledge, and are prone to making prediction errors.
9. Conclusion
In this paper, we introduced SmartHalo, a novel framework designed to address critical deficiencies in existing Solidity smart contract decompilers. By leveraging the strengths of static analysis (SA) and large language models (LLM), SmartHalo significantly enhances the accuracy and reliability of decompiled code. Our framework systematically recovers method boundaries, variable types, and contract attributes, facilitating the program comprehension and vulnerability detection. Extensive evaluation demonstrates that SmartHalo outperforms state-of-the-art decompilers, showing substantial improvements in method boundary identification, variable type inference, and contract attribute recovery. We also showed that the output of SmartHalo can significantly enhance the effectiveness of downstream tasks for vulnerability detection.
Data Availability
Currently, the artifact of SmartHalo, corresponding evaluation dataset, and other online materials are available in this anonymous repository (https://figshare.com/s/5d4fc1ad2312a4be9370).